# Operating System And Artificial Intelligence: A Systematic Review

### v1

@@ 6.1 Tools analysis
A decision tree is a type of machine learning model used when the relationship between a set of predictor variables and a response variable is non-linear,while random forest is essentially a collection of decision trees. It is quick to fit to a dataset and easy to interpret. Chowdhury et al. [32] used RF as one if the models to accurately detect the attack from the network traffic. To construct the random forest classifier, the authors employed the Random Forest Regressor , which serves as both a regressor and a meta estimator. It accomplishes this by fitting multiple decision trees to different subsets of the dataset. For our specific model, we opted for a forest containing 1000 trees.Ahmed et al. [23] used RF to build a device fitness model,based on the Dataset collected during runtime and statically. De Wit et al.[39] trained a statistical classifier able to recognize malware signatures in any log data collected on a smartphone. The classifier was trained, cross-validated, and tested using the dataset described above and RF classifier had a better performance.Metzge el at.[49] used RF to get a optimal kernel runtime switching slice size. The model is a random forest regressor with 50 decision trees with a depth of two for the GPU model. Ongun et al.[36] used RF to get the probability of a command being malicious,based on labels dataset.
@@
RNNs are a class of neural networks designed to handle sequential data. They have feedback connections, allowing them to maintain an internal state or memory. Each step in an RNN processes an input and updates its hidden state based on the current input and the previous hidden state.While LSTM ,a type of RNN architecture with a more complex cell structure,were introduced to address the vanishing gradient problem. Motivated by the problem that exploiting 3D-stacking memory’s performance is challenging because bandwidth utilization heavily depends on address mapping in the memory controller, Zhang et al.[17] used a software-defined address mapping, allowing user programs to directly control low-level memory hardware in an intelligent and fine-grained manner.LSTM is used to in a method to get access pattern information to select an address mapping. It identify the major variables that significantly contribute to external memory access and have a substantial impact on memory traffic and data movement.
WordEmbedding
Word embeddings provide a way to achieve this by mapping words to dense vectors in a multi-dimensional space. Fu et al.[34] proposed a AI-based approach to help under-resourced security analysts to find, detect,and localize vulnerabilities. They utilized a word-level Clang tokenizer with a copy mechanism. This tokenizer broke down a C function into a sequence of tokens.Then word embedding was used to generate vector representations for each token in the sequence, capturing the semantic information among the input tokens. And further classifier was , Ongun et al.[36] explore techniques for representing command-line text using word embedding methods.Based on this,they devise ensemble boosting classifiers to differentiate between malicious and non-malicious commands.
@@
KNN is a supervised machine learning algorithm used for both classification and regression tasks. The fundamental idea behind KNN is simple: neighbors influence each other. If you’re surrounded by similar things, you’re likely similar too. It is widely applicable in pattern recognition, data mining, and intrusion detection. Yang et al[42] introduce a KNN-based machine learning algorithms can accurately predict the Turnaround-time(TaT) of a process. It can effectively reduce the TaT of the process and reduce the number of process context switches.
@@
An MLP is a type of feedforward neural network used for supervised learning tasks, such as classification and regression. Chen et al.[3] argues that traditional Linux CFS scheduler maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources and try to solve the above problem using an ML-based resource-aware load balancer.They employed supervised imitation learning to replace a portion of its internal logic with an MLP model. This trained MLP model emulates the kernel’s load balancing decisions. MLP is chosen because this current work doesn’t require a very complex model and MLP has a relatively simple implementation compared to the other models. Based on this work,Qiu et al.[78] propose the concept of reconfigurable kernel datapaths that enables kernels to self-optimize dynamically to reduce the cost of kernel. The authors also used MLP ML model that can mimic Linux CFS decisions.
6.2 Findings
In addressing Research Question 3, which explores the methods by which AI techniques are utilized to enhance OS, our analysis reveals a variety of approaches tailored to meet distinct OS requirements. These AI techniques are strategically chosen to address a spectrum of OS challenges, encompassing security enhancements, performance forecasting, resource management, and process scheduling. The widespread adoption of these methods attests to their versatility and underscores the profound influence of AI in augmenting OS functionality and performance. This integration of AI into core OS components not only boosts operational efficiency but also paves the way for more intelligent and adaptable systems, capable of meeting the evolving demands of modern computing environments.
7 RQ4: How and what novel OS architectures create opportunities for AI4OS?
@@
7.1 Kernel-bypass OS Structure for AI
Prior research has successfully demonstrated the potential of leveraging hardware acceleration for machine learning within the kernel space, showcasing the feasibility of such an approach[4]. This study introducES an API remoting system, which facilitates access to specialized accelerator interfaces for kernel space applications. Moreover, it simplifies the integration by offering high-level APIs directly to the kernel space, eliminating the necessity for kernel-specific adaptations of complex libraries. The API remoting mechanism transmits commands between kernel and user space. This innovative design not only enhances the performance of AI applications within the kernel but also provides a compelling perspective on the capabilities of modern operating systems.
Raza et al. [53] propose integrating unikernel optimizations into Linux, known for creating secure, compact OS images for single applications. Unikernel Linux (UKL) reduces the number of executed instructions and improves instructions-per-cycle efficiency. Tail latency tests on Memcached—a multi-threaded key-value store—show that UKL achieves a significant performance improvement. It introduces a configuration option that allows a single, optimized process to link directly with the kernel, bypassing the traditional system call overhead, and significantly cuts latency for system call payloads, showcasing the benefits of unikernel-inspired enhancements in refining Linux’s performance.
Cadden et al.[59] introduces Serverless Execution via Unikernel SnapShots (SEUSS), a method that leverages unikernel snapshots for rapid deployment and high-density caching of serverless functions. The authors describe the use of unikernel contexts, which consist of a high-level language interpreter configured to import and execute function code, providing isolation and security. This minimalistic approach leads to a reduced memory footprint and faster startup times compared to traditional operating systems like Linux, which is beneficial for serverless environments where rapid function instantiation is crucial. By using unikernels, the SEUSS system can cache a large number of function instances in memory due to the reduced memory footprint.
@@ 7.2 Library OS
The Demikernel project[52] unveils an OS architecture optimized for datacenter systems with microsecond-scale requirements, emphasizing low-latency I/O. Adopting a LibOS strategy, it side-steps the traditional kernel in I/O paths, significantly enhancing performance. This design supports kernel-bypass devices, allowing seamless application operation with negligible overhead at the nanosecond scale. For I/O efficiency, Demikernel applies zero-copy techniques for large buffers, optimizes resource use, and maintains system stability through periodic LibOS interaction. Security is bolstered with controlled data placement and potential advanced memory integrity measures. LibOS components employ hardware acceleration to offload critical tasks, minimizing latency and maximizing throughput. Kernel bypass in I/O paths reduces overhead from context switching, system calls, and memory duplication.
@@ 7.3 Other Stuctures
Skiadopoulos et al. [60] present DBOS as a superior alternative to conventional cluster OS components, offering comparable functionality but enhanced analytics and reduced code complexity. DBOS matches current systems in scheduling, file management, and inter-process communication, yet excels in analytics and simplifies code through database query-based OS services. It efficiently implements low-latency transactions and ensures high availability. DBOS’s integrated DBMS approach is especially advantageous for ML, delivering a cohesive platform for efficient resource management and analytics in large-scale distributed environments, adeptly managing parallel computation and workload dynamics across various hardware.
@@
Shan et al.[61] developed LegoOS—a splitkernel architecture for hardware disaggregation, decentralizing OS functions for scalable, distributed management. It fundamentally breaks down traditional OS functionalities into loosely-coupled monitors that each run on and manage a distinct hardware component. The splitkernel model distributes responsibilities such as scheduling, memory management, and I/O operations across these monitors, effectively creating a distributed set of hardware components that can be independently managed and scaled. The splitkernel model in LegoOS allows for independent scaling of compute and memory resources, essential for ML scenarios where large datasets and complex models necessitate specialized hardware accelerators.
7.4 Findings
Addressing Research Question 4, our analysis highlights the transformative role of novel OS architectures in advancing AI4OS. Pioneering designs like LibOS and unikernels have sparked renewed excitement among researchers, demonstrating their potential to outshine traditional OSes. These cutting-edge architectures enhance the symbiosis between software and hardware, reducing latency to boost AI application responsiveness and real-time processing capabilities. Their streamlined construction optimizes resource allocation, ensuring efficient management of computational assets. Beyond performance, these architectures establish a secure foundation for AI deployment. By simplifying the environment, they minimize security risks, vital for AI applications managing sensitive data. DBOS and LegoOS exemplify this by simplifying distributed AI setup and maintenance, freeing experts to focus on algorithmic innovation. They also adapt nimbly to the variable demands and hardware diversity of modern ML, enhancing deployment efficiency.

### v2

@@ 4.7.2 How to choose proper algorithms to improve OS?
A decision tree is a type of ML model used when the relationship between a set of predictor variables and a response variable is non-linear,while random forest is essentially a collection of decision trees. It is quick to fit to a dataset and easy to interpret. Chowdhury et al. (?) used RF as one if the models to accurately detect the attack from the network traffic. To construct the random forest classifier, the authors employed the Random Forest Regressor, which serves as both a regressor and a meta estimator. It accomplishes this by fitting multiple decision trees to different subsets of the dataset. For this specific model, the authors opted for a forest containing 1000 trees. Ahmed et al. (?) used RF to build a device fitness model,based on the Dataset collected during runtime and statically. De Wit et al. (?) trained a statistical classifier able to recognize malware signatures in any log data collected on a smartphone. The classifier was trained, cross-validated, and tested using the dataset described above and RF classifier had a better performance. Metzge el at. (?) used RF to get a optimal kernel runtime switching slice size. The model is a random forest regressor with 50 decision trees with a depth of two for the GPU model. Ongun et al. (?) used RF to get the probability of a command being malicious,based on labels dataset. Sun et al. (?) used RF to filter storage I/O parameters demonstrates that tuning only the important parameters can nearly achieve the high performance obtained by tuning all parameters. Known for their versatility in handling non-linear relationships and minimal tuning requirements, RF/DT are used for tasks such as attack detection and configuration settings. They are chosen for their ability to quickly fit to datasets and provide interpretable models.
@@
RNNs are a class of neural networks designed to handle sequential data. They have feedback connections, allowing them to maintain an internal state or memory. Each step in an RNN processes an input and updates its hidden state based on the current input and the previous hidden state. While LSTM, a type of RNN architecture with a more complex cell structure,were introduced to address the vanishing gradient problem. Motivated by the problem that exploiting 3D-stacking memory’s performance is challenging because bandwidth utilization heavily depends on address mapping in the memory controller, Zhang et al. (?) used a software-defined address mapping, allowing user programs to directly control low-level memory hardware in an intelligent and fine-grained manner. LSTM is used to in a method to get access pattern information to select an address mapping. It identify the major variables that significantly contribute to external memory access and have a substantial impact on memory traffic and data movement. Miura et al. (?) proposed to apply LSTM for time series data, for cache replacement in the second cache. Designed to handle sequential data, RNN/LSTM are particularly suited for tasks that require capturing temporal dependencies, such as managing memory access patterns in 3D-stacking memory systems.
@@
KNN is a supervised ML algorithm used for both classification and regression tasks. The fundamental idea behind KNN is simple: neighbors influence each other. If you’re surrounded by similar things, you’re likely similar too. It is widely applicable in pattern recognition, data mining, and intrusion detection. Yang et al. (?) introduce a KNN-based ML algorithms can accurately predict the Turnaround-time (TaT) of a process. It can effectively reduce the TaT of the process and reduce the number of process context switches.
@@
An MLP is a type of feedforward neural network used for supervised learning tasks, such as classification and regression. Chen et al. (?) argues that traditional Linux CFS scheduler maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources and try to solve the above problem using an ML-based resource-aware load balancer.They employed supervised imitation learning to replace a portion of its internal logic with an MLP model. This trained MLP model emulates the kernel’s load balancing decisions. MLP is chosen because this current work doesn’t require a very complex model and MLP has a relatively simple implementation compared to the other models. Based on this work,Qiu et al. (?) propose the concept of reconfigurable kernel datapaths that enables kernels to self-optimize dynamically to reduce the cost of kernel. The authors also used MLP ML model that can mimic Linux CFS decisions.
5 Large Language Model Enhanced Operating Systems
The rapid development of LLMs has introduced transformative capabilities in natural language understanding and generation, making them a promising tool for advancing operating system (OS) design and interaction. LLMs, such as GPT-4 (?), LLaMA (?), and Gemini (?), have demonstrated remarkable proficiency in processing and generating human-like text, understanding complex instructions, and adapting to a wide variety of tasks with minimal or no task-specific training. These models are built on advances in deep learning architectures, particularly Transformer-based designs, and trained on vast datasets that enable them to capture context, semantics, and nuanced relationships between concepts.
One of the defining features of LLMs is their ability to act as general-purpose reasoning and interaction tools. Unlike traditional machine learning models, which are often task-specific, LLMs are highly adaptable, capable of handling a broad range of tasks through natural language interfaces. Their ability to understand intent, provide detailed explanations, and generate contextualized responses makes them an ideal complement to OS functionalities. As operating systems serve as the backbone for user-computer interaction, integrating LLMs can enable more intuitive, conversational interfaces, where users can communicate their needs and preferences in natural language rather than relying on rigid command structures or graphical user interfaces. The integration of LLMs into OSs presents a significant opportunity to enhance the user experience and overall system functionality (?, ?). LLMs can transform the way users interact with and manage their computing environments (?).
5.1 Development
LLMs possess powerful natural language processing capabilities, making them highly effective at understanding historical documentation on human experiences with operating systems. This ability is particularly valuable in the era of the Internet of Everything, where engineers invest significant effort in adapting hardware and applications to achieve optimal resource utilization and efficiency.
Kamath et al.’s LLaMaS (?) and Chen et al.’s AutoOS (?) both utilize LLMs to address challenges in operating system configuration optimization, but they focus on different domains and employ distinct methodologies. LLaMaS addresses the complexities of adapting to diverse hardware environments by using LLMs to interpret plain text hardware specifications and recommend optimized OS strategies. Its architecture comprises a frontend and a backend: the frontend deciphers system characteristics and converts them into actionable embeddings, which the backend then uses for real-time OS decisions, such as managing memory tier data movement. In an experiment, LLaMaS demonstrated its capability by using ChatGPT to dynamically adjust memory allocation for CPU and GPU tasks based on workload demands. By leveraging LLMs’ zero-shot learning capabilities, LLaMaS eliminates the need for manual adjustments or device-specific coding, significantly reducing administrative and development costs. In contrast, AutoOS tackles the challenge of Linux kernel configuration optimization, a critical task for AIoT devices with constrained hardware resources. AutoOS models the kernel configuration space as a dynamic tree structure and employs a state-machine-based framework to traverse, prune, propose, and correct configuration options. The LLM frontend identifies performance-critical options, while the backend applies these refinements to boost system performance. Experiments showed that AutoOS could enhance UnixBench scores by up to 25% compared to vendor-provided defaults, achieving superior results with minimal human input. By automating the debugging and refinement of kernel configurations, AutoOS reduces the time and expertise required for manual optimization, addressing the complexity of configuring over 15,000 kernel options.
Both LLaMaS and AutoOS leverage LLMs to simplify traditionally complex OS tasks, minimizing the need for manual tuning and expert knowledge. They also demonstrate the ability of LLMs to adapt dynamically to changing requirements. Moreover, both frameworks utilize LLMs’ reasoning and zero-shot learning capabilities to autonomously make decisions, showcasing their effectiveness in reducing development and administrative overhead.
Another important and obvious area where LLMs excel is coding. As a type of textual data, code can be considered another form of human language, making LLMs particularly well-suited for tasks involving programming. This highlights the crucial role LLMs can play in improving software development processes, from code generation to debugging and vulnerability repair. For instance, SecRepair (?) is a system specifically designed to address the challenge of identifying and repairing code vulnerabilities. Powered by an LLM, it integrates reinforcement learning and a semantic reward mechanism to enhance its performance. To support its training, the authors have compiled and released a comprehensive instruction-based vulnerability dataset, which serves as a key resource for building a robust model for vulnerability analysis. Furthermore, SecRepair employs a reinforcement learning technique with semantic rewards to generate concise and contextually appropriate code commit comments. This approach, inspired by how humans resolve code issues, not only automates the repair process but also provides developers with a clear understanding of the vulnerabilities and their resolutions, ultimately streamlining software development and improving code security.
Insights and Challenges:
LLMs have the potential to significantly enhance software development processes, including tasks such as code generation, debugging, and vulnerability detection, by reducing manual effort and improving overall efficiency. The development and release of comprehensive, instruction-based datasets are essential for training models capable of addressing complex tasks like vulnerability analysis and repair. However, several challenges remain. For instance, mainstream operating systems like Linux consist of vast codebases with deeply coupled structures, making it difficult for current LLMs to fully capture all necessary modifications when proposing adaptations or optimizations. While LLMs can provide efficient suggestions, they may overlook related areas in the code that also require changes. Additionally, training a new LLM model demands significant time and computational resources, posing challenges in scenarios where engineers frequently switch to new hardware platforms. It is often impractical and uneconomical to train a fresh model for each new hardware adaptation, highlighting the need for more efficient and scalable solutions.
5.2 User Interaction
LLMs are transforming user interaction by enabling more natural, intelligent, and personalized experiences across various systems. As the critical bridge between hardware and software, the operating system plays a pivotal role in shaping the user interaction experience, which in turn defines how users engage with and perceive technology. Researchers are now striving to bring this revolutionary shift into the realm of operating systems, aiming to redefine how users interact with computing environments.
Personalized services can meet specific user needs, enhancing satisfaction and loyalty. Furthermore, interactive systems can understand users’ perspectives and needs through well-designed interfaces and experiences. Zhang et al. (?) explores the potential applications of large language models, ML, and interaction design in recommendation systems and operating systems. By integrating these technologies, more intelligent and personalized services can be provided, meeting user needs and promoting continuous improvement and optimization of products, which holds great value for both recommendation research and user experience applications.
Some researchers are exploring deeper structural reforms in OSs through the integration of LLMs. The AIOS-Agent ecosystem (?, ?) positions large language models (LLMs) as an ”operating system with a soul,” embedding them into the OS kernel to enable intelligent decision-making and resource management in Figure.6. The LLM’s context window functions as memory, handling relevant data, while external storage acts as a file system with advanced retrieval capabilities. Hardware and software are treated as peripherals and libraries, facilitating seamless agent-environment interaction. Natural language serves as the primary programming interface, democratizing software development. This ecosystem supports both single- and multi-agent applications capable of executing a wide range of tasks. Building on the AIOS concept, various studies have further enriched the ecosystem with additional components.
Figure 5: Architecture for AIOS
Figure 6: Architecture for LSFS
Expanding on this foundation, the CoRE (Code Representation and Execution) system (?) introduces a unified framework for natural language programming, pseudo-code programming, and flow programming. CoRE leverages LLMs as interpreters to execute logically structured programs described in natural language, bridging the gap between human intent and machine execution. By defining a structured syntax, the CoRE language allows users to articulate program logic in a manner that is both accessible to non-programmers and precise enough for computational interpretation. Programs are decomposed into components such as step names, types, instructions, and connections, enabling clear and logical execution.
Within the AIOS ecosystem, the LLM-based Semantic File System (LSFS) (?) in figure.6 redefines traditional file management by integrating LLMs for natural language-driven interactions and semantic understanding of file content. Unlike conventional systems that rely on metadata such as file names and timestamps, LSFS employs embedding vectors to index files based on their content, enabling advanced functionalities like semantic retrieval, grouping, and version control. Users can perform complex operations—such as file updates and change summarization—through simple natural language prompts, significantly streamlining file management. Key features include semantic indexing, natural language parsing for file operations, real-time change monitoring, and robust safety mechanisms to prevent unintended actions. Seamlessly integrated into the AIOS ecosystem, LSFS enhances intelligent multi-agent collaboration and supports capabilities such as task scheduling and knowledge integration. This innovation transforms file systems into intelligent, user-friendly components that are essential for modern AI-driven environments.
Insights and Challenges:
LLMs are driving a paradigm shift in user interaction by enabling natural, intuitive, and personalized experiences. Their integration into OSs redefines how users interact with hardware and software, making technology more accessible and user-centric. Users can interact with computing environments in a more intuitive manner, leading to increased satisfaction and loyalty. The shift towards personalized services allows systems to adapt to specific user needs, leveraging insights from interaction data to continuously refine and optimize user experiences. For instance, the work of researchers like Zhang et al. (?) highlights how LLMs can enhance recommendation systems. This trend underscores the importance of understanding users’ perspectives, enabling systems to respond dynamically to their needs through well-crafted interfaces. Moreover, innovative frameworks such as AIOS-Agent and CoRE exemplify the potential of LLMs in redefining OS functionality. By embedding LLMs into the OS kernel and employing natural language as a programming interface, these systems democratize software development and empower users—regardless of technical background—to engage with technology more effectively.
Despite these promising developments, several challenges remain in the integration of LLMs within operating systems. One significant concern is the lack of interpretability associated with LLMs, which can impact their reliability. Users and developers may struggle to understand how LLMs arrive at specific outputs, leading to a lack of trust in their decisions and recommendations. This opacity can complicate troubleshooting and diminish user confidence, particularly in critical applications where understanding the rationale behind actions is essential. Additionally, there are challenges related to the scalability of these systems. As LLMs require significant computational resources, optimizing their performance in diverse hardware environments remains a critical hurdle. This is particularly important in ensuring that systems can deliver real-time responses without compromising efficiency. Moreover, while LLMs can enhance user interaction, they also raise questions about privacy and data security. The reliance on user data to personalize experiences necessitates robust safeguards to protect sensitive information and maintain user confidentiality.
5.3 Analysis
Existing research has explored various ways in which LLMs can be leveraged within OS architectures. LLaMaS, for instance, addresses the complexities of diverse computing environments by using LLMs to ease OS challenges in hardware adaptation and resource management. AIOS-Agent ecosystem treats LLMs as an ”operating system with soul”, embedding them in the OS kernel for intelligent decision-making and resource allocation. MemGPT introduces virtual context management to overcome the fixed-length context window limitation in LLMs, enabling them to access information beyond their immediate capacity. Additionally, systems like SecRepair and ChronoCTI demonstrate how LLMs can be employed to address challenges in software vulnerability repair and cybersecurity threat analysis, respectively.
5.3.1 Differnce compared to ML-enhanced OS:
LLMs and traditional ML techniques enhance OS in distinct ways, reflecting their unique capabilities and methodologies.
Traditional ML techniques are often more task-specific and rely on structured data for training. These methods are typically employed for optimization tasks, such as system performance tuning, resource allocation, and configuration management. Additionally, ML models are frequently viewed as black boxes, making it challenging for users to understand the underlying principles and the relationship between the input training data and the resulting outputs. This lack of interpretability can hinder trust and usability, especially in critical applications where comprehending the rationale behind decisions is essential. In contrast, LLMs primarily focus on natural language understanding and generation, enabling more intuitive user interactions through conversational interfaces. They excel in interpreting user intent, managing complex tasks through natural language commands, and providing personalized experiences based on user input. This integration allows LLMs to transform how users engage with the OS, making technology more accessible and user-friendly.
While there has been extensive research exploring the integration of traditional ML techniques into operating systems, where they effectively delve into the complexities of system performance, resource allocation, and configuration management, the role of LLMs remains more superficial. LLMs primarily function as an additional layer that enhances user interaction with the OS through natural language processing, enabling more intuitive and conversational interfaces.
6 How do novel OS architectures create opportunities for intelligent OS?
@@
6.1 Kernel-bypass OS Structure
Prior research has successfully demonstrated the potential of leveraging hardware acceleration for ML within the kernel space, showcasing the feasibility of such an approach (?). This study introducES an API remoting system, which facilitates access to specialized accelerator interfaces for kernel space applications. Moreover, it simplifies the integration by offering high-level APIs directly to the kernel space, eliminating the necessity for kernel-specific adaptations of complex libraries. The API remoting mechanism transmits commands between kernel and user space. This innovative design not only enhances the performance of AI applications within the kernel but also provides a compelling perspective on the capabilities of modern operating systems.
Raza et al. (?) propose integrating unikernel optimizations into Linux, known for creating secure, compact OS images for single applications. Unikernel Linux (UKL) reduces the number of executed instructions and improves instructions-per-cycle efficiency. Tail latency tests on Memcached—a multi-threaded key-value store—show that UKL achieves a significant performance improvement. It introduces a configuration option that allows a single, optimized process to link directly with the kernel, bypassing the traditional system call overhead, and significantly cuts latency for system call payloads, showcasing the benefits of unikernel-inspired enhancements in refining Linux’s performance.
Cadden et al. (?) introduces Serverless Execution via Unikernel SnapShots (SEUSS), a method that leverages unikernel snapshots for rapid deployment and high-density caching of serverless functions. The authors describe the use of unikernel contexts, which consist of a high-level language interpreter configured to import and execute function code, providing isolation and security. This minimalistic approach leads to a reduced memory footprint and faster startup times compared to traditional operating systems like Linux, which is beneficial for serverless environments where rapid function instantiation is crucial. By using unikernels, the SEUSS system can cache a large number of function instances in memory due to the reduced memory footprint.
@@ 6.2 Library OS
The Demikernel project (?) unveils an OS architecture optimized for datacenter systems with microsecond-scale requirements, emphasizing low-latency I/O. Adopting a LibOS strategy, it side-steps the traditional kernel in I/O paths, significantly enhancing performance. This design supports kernel-bypass devices, allowing seamless application operation with negligible overhead at the nanosecond scale. For I/O efficiency, Demikernel applies zero-copy techniques for large buffers, optimizes resource use, and maintains system stability through periodic LibOS interaction. Security is bolstered with controlled data placement and potential advanced memory integrity measures. LibOS components employ hardware acceleration to offload critical tasks, minimizing latency and maximizing throughput. Kernel bypass in I/O paths reduces overhead from context switching, system calls, and memory duplication.
@@ 6.3 Other Stuctures
Skiadopoulos et al. (?) present DBOS as a superior alternative to conventional cluster OS components, offering comparable functionality but enhanced analytics and reduced code complexity. DBOS matches current systems in scheduling, file management, and inter-process communication, yet excels in analytics and simplifies code through database query-based OS services. It efficiently implements low-latency transactions and ensures high availability. DBOS’s integrated DBMS approach is especially advantageous for ML, delivering a cohesive platform for efficient resource management and analytics in large-scale distributed environments, adeptly managing parallel computation and workload dynamics across various hardware.
@@
Shan et al. (?) developed LegoOS—a splitkernel architecture for hardware disaggregation, decentralizing OS functions for scalable, distributed management. It fundamentally breaks down traditional OS functionalities into loosely-coupled monitors that each run on and manage a distinct hardware component. The splitkernel model distributes responsibilities such as scheduling, memory management, and I/O operations across these monitors, effectively creating a distributed set of hardware components that can be independently managed and scaled. The splitkernel model in LegoOS allows for independent scaling of compute and memory resources, essential for ML scenarios where large datasets and complex models necessitate specialized hardware accelerators.
6.4 Analysis
All these works agree that, facing diverse computing environments, traditional OSs have encountered significant challenges. They reflect the increasing importance of AI in computing and the need for OS architectures that can effectively support AI workloads. They all strive to adapt to Heterogeneous Devices and Diverse Computing Needs. They focus on transforming low-level structures for efficiency and performance optimization in different aspects. Unikernels aim to enhance performance by reducing overhead and streamlining interactions between software and hardware. LibOS and split-kernels primarily focus on simplified integration and streamlined architectures, emphasizing scalability and flexibility. These works together pave a solid foundation for AI to refactor the structures of OSs, enabling better alignment with modern computing demands.
Our analysis reveals several key innovations in OS architecture that offer new possibilities for the integration and application of AI:
Kernel-Bypass OS Structures:
Researchers have demonstrated the potential of leveraging hardware acceleration for machine learning within the kernel space. For instance, an API remoting system allows kernel space applications to directly access specialized accelerator interfaces, enhancing the performance of AI applications.
Library OS:
Addressing the challenges of datacenter systems requiring microsecond-scale responses, the Demikernel project introduces an OS architecture optimized for such demands. This LibOS strategy bypasses traditional kernels in I/O paths, significantly improving performance. LibOS components utilize hardware acceleration to offload critical tasks, minimizing latency and maximizing throughput.
Other Structures:
DBOS presents itself as a superior alternative to conventional cluster OS components, offering comparable functionality with enhanced analytics and reduced code complexity. It efficiently implements low-latency transactions and ensures high availability, making it especially advantageous for ML in large-scale distributed environments. LegoOS, developed as a split-kernel architecture for hardware disaggregation, decentralizes OS functions for scalable, distributed management, allowing independent scaling of compute and memory resources.
These innovative OS architectures not only enhance performance and security but also provide better support for the needs of AI applications. They ensure efficient management through simplified environments and optimized resource allocation, providing a secure foundation for AI deployment. DBOS and LegoOS, by simplifying distributed AI setup and maintenance, allow experts to focus on algorithmic innovation. These architectures adapt nimbly to the variable demands and hardware diversity of modern ML, enhancing deployment efficiency.